OpenStudy (bobobox):

A.Create and justify a relation that represents a function with the given range. Write your answers as ordered pairs. B.Create and justify a relation that does not represent a function with the given range. Write your answers as ordered pairs.

OpenStudy (texaschic101):

whats the given range ?

OpenStudy (bobobox):

I'm working on attaching it uno momento

OpenStudy (texaschic101):

let me try and explain.. the range is all the y values and the domain is all the x values it is a function if there is only 1 x value for every y value. There cannot be 2 or more x's the same, but there can be 2 or more y's the same. So your range (y values) are 11,20,34,41 your x values can be any number just as long as they are not repeating x numbers. For example : (1,11)(2,20)(3,34)(4,41) --- this is a function because there is no repeating x values. (1,11)(1,20)(2,34)(4,41)--- this is not a function because it has 2 x values that are the same. does this make any sense to you ??

OpenStudy (bobobox):

I have been through this lesson forward and backwards, I get the concept but not really the question where I live it is 11:44 I'm tired this was due a week ago help a brother out.

OpenStudy (texaschic101):

lol....I will help a brother out.....answer is : 1,2,3,4...these are your x values

OpenStudy (texaschic101):

its only 9:44 where I live

OpenStudy (bobobox):

thank you so very much, not to sound greedy/needy/lazy (hehe) could you help me with one more? I will fan and write and awesome testimony...

OpenStudy (texaschic101):

I can try

OpenStudy (bobobox):

Find the equation of the line in the graph, and mathematically model the scenario using function notation where cost is a function of number of minutes.

OpenStudy (texaschic101):

ask away

OpenStudy (bobobox):

wait for it...

OpenStudy (texaschic101):

lets pick 2 points...(200,30) and (400,40) find the slope using slope formula : slope=(y2 - y1) / (x2 - x1) slope(m) = (40 - 30) / (400 - 200) slope(m) = 10/200 slope(m) = 1/20 now use y = mx + b slope(m) = 1/20 use either of your points (200,30)...x= 200 and y = 30 now sub 30 = 1/20(200) + b 30 = 200/20 + b 30 = 10 + b 30 - 10 = b 20 = b equation is : y = 1/20x + 20 function notation is : f(x) = 1/20x + 20....I think thats how you do it, but you might eed a second opinion Lets check.. f(x) = 1/20 + 20 f(400) = 1/20(400) + 20 f(400) = 20 + 20 f(400) = 40 <=== here 400 minutes cost $40, just like your graph

OpenStudy (texaschic101):

I am sorry but I have to go now...go get some rest...lol

OpenStudy (bobobox):

Holy crapola you deserve an awesome testimony and a hug from the algebra go himself...Thank you so much!